# Direct numerical simulations of a cylinder cutting a vortex

### v1

@@ III.1 The weak and strong vortex regimes
Turning to the weak vortex regime, Figure 7 shows volume plots of the vorticity magnitude $|\omega|$ for $Re_{\Gamma}=1000$ and $I_{P}=0.25$. As in the previous case, two sources of vorticity are evident: the primary vortex and the wire’s boundary layer. However, the overall flow evolution looks very different when compared to Figure 5. The presence of a weaker vortex results in lower induced velocities on the wire. The wire’s boundary layer fails to lift off, so there is no secondary vorticity curling around and deforming the primary vortex before impact. Consequently, the wire’s boundary layer does not play a significant role in this scenario. This lack of interaction before the wire approaches the core serves as a key characteristic of the weak vortex regime, and has facilitated its modeling using inviscid methods.
@@
Instead, in this regime, the core deformation and the propagation of waves dominate, as illustrated in the second and third panels. These deformations arise from the physical presence of the wire traversing the vortex tube, which displaces it and stretches it until the core is eventually “cut”. The point-wise deformation propagates through waves along the primary tube’s axis, previously observed in Figure 3. These do not substantially modify the radius of the primary vortex, a phenomena observed in the simulations and experiments of weak-vortex BVI with axial flow in Ref. [22]. This flow behavior differs from BVI with axial flow in the weak vortex regime. Axial flow introduces a notable asymmetry in wave propagation, an aspect absent in our current observations. Additionally, the object’s presence obstructs the axial flow, creating “compression” and ’expansion’ regions that modify the vortex radius [22].
Finally, the fourth panel depicts the wire after cutting through the tube, leaving a wake that interacts with the strongly deformed vortex tube. Over time, the vortex manages to reconstitute itself and survive the cutting, albeit with a significantly altered vorticity profile, as shown in the last panel. The primary vortex has also entrained a small quantity of vorticity from the boundary layer, as suggested by Ref. [22]. We also note that in this case, we were able to observe some of the phenomena usually attributed to the bending regime, i.e. deformations of the primary vortex, and a stretching of the core near the impact location. This hints at the fact that the transition between strong vortex and bending regimes, which is outside the scope of this paper, will also be smooth.
@@
Figure 7: Visualization of the time evolution of $|\omega|$ for a simulation in the weak vortex regime with $Re_{\Gamma}=1000$, $I_{P}=0.25$. Snapshots are at $t\Gamma/D^{2}=200$, $240$, $280$, $320$ and $440$.
@@
These observations are underscored in Figure 8, where the vorticity at the center $x$-$z$ plane is depicted for four distinct time points—prior to, during, and after the core cutting. In the first panel, we show the wire approaching the vortex, which remains as a circular path of vorticity. In the second panel, the vorticity originating from the wire is beginning to wrap around the vortex, but at a slow rate. It is not fast enough to surround the vortex before the main wire reaches the core. Notably, although the vorticity has less time to detach, it exhibits a higher value than in Figure 6. This can be attributed to the non-dimensionalization using $\Gamma/\sigma$ as a velocity scale, where higher values of $I_{P}$ lead to greater wire velocities, resulting in larger vorticity values for a boundary layer when normalized using $\Gamma$-$\sigma$ units.
In the third panel, the vortex has already shifted from its initial position, with intensified vorticity due to stretching. Finally, in the fourth panel, the wire has completely traversed the vortex, leaving behind vorticity associated with the primary tube. The distribution differs, with vorticity concentrated in a thin sheet and displaced to the right by a few vortex cores. Despite this, a comparison of the initial and final circulation of the vortex reveals only a $5\%$ loss throughout the entire process. This showcases the vortex’s ability to withstand the interaction significantly better than the case at $I_{P}=0.05$, which suffered a $75\%$ loss of its original circulation.
@@
Figure 8: Visualization of the time evolution of $|\omega|$ for a simulation in the weak vortex regime with $Re_{\Gamma}=1000$, $I_{P}=0.25$. Snapshots are at $t\Gamma/D^{2}=120$, $200$, $240$ and $400$.
@@
We can quantify the loss of circulation after the interaction has taken place ($\Gamma_{f}$) for all cases. This is shown in figure 9, where we plot $\Gamma_{f}/\Gamma_{0}$, $\Gamma_{0}$ being the original circulation of the vortex. We observe from the plot that the amount of circulation in the vortex by the end of the interaction monotonically increases with $I_{P}$ in the range studied. We note that this trend cannot persist, as the circulation must saturate at $\Gamma_{f}\approx\Gamma_{0}$ at higher values of $I_{P}$, as it is not expected that the wire imparts additional circulation to the vortex.
@@
III.2 The effect of Reynolds number
Up to now our discussion has been limited to $Re_{\Gamma}=1000$. Many hydrodynamical instabilities of vortex tubes set in at values of $Re_{\Gamma}$ which are similar or slightly higher [36]. Therefore, expanding our analysis to higher Reynolds numbers will not only assess the Reynolds number’s influence on BVI but could also lead to very different outcomes due to the heightened instability of the resulting vortices.
To explore this, we commence by visualizing the time evolution of vorticity in the strong vortex regime at $Re_{\Gamma}=3000$ and $I_{P}=0.05$ in Figure 10, which offers a comparison with the $Re_{\Gamma}=1000$ counterpart shown in Figure 5. An initial inspection reveals seemingly minor differences between both cases. The lifted secondary vorticity from the boundary layer exhibits higher values but looks thinner, consistent with the increased Reynolds number of the wire. The fundamental flow mechanisms seem largely unaffected: a thinning of the core near the impact point, the formation of a neck-like structure, and an increase of vorticity. Moreover, the secondary vortices lifting off and wrapping around the vortex propagate through its axis at a comparable speed. Key disparities include heightened distortion of the primary vortex, now featuring smaller wavelength components, and that the turbulent cloud left behind is formed from smaller structures. Notably, unlike the $Re_{\Gamma}=1000$ scenario where remnants persisted, at $Re_{\Gamma}=3000$, we cannot identify a similar remnant of the primary vortex after the interaction has taken place, preventing the calculation of a value for $\Gamma_{f}$ as previously done.
Figure 10: Visualization of the time evolution of $|\omega|$ for a simulation in the strong vortex regime with $Re_{\Gamma}=3000$, $I_{P}=0.05$. Snapshots are at $t\Gamma/D^{2}=740$, $1140$, $1540$ and $1740$.
To better appreciate the effect of the higher Reynolds number, we show the vorticity magnitude in a cut through the $z=0$ plane just before the impact of the cylinder for the two Reynolds numbers simulated. The flow features remain qualitatively similar, even if the right panel shows sharper features: the wrapping of vorticity around the primary tube, as well as its role in forming the neck can be observed. This underscores the importance of secondary vorticity in the strong vortex regime, as postulated in Ref. [17]. Models of normal BVI within the strong vortex regime must adequately incorporate this phenomenon as it plays a pivotal role in driving the dynamics.
Figure 11: Vorticity magnitude in the $z=0$ plane just before impact ($t\Gamma/D^{2}=940$) for $I_{P}=0.05$ and $Re_{\Gamma}=1000$ (left) and $Re_{\Gamma}=3000$ (right).
Turning to the weak vortex case ($I_{P}=0.25$), we present visualizations of the vorticity modulus in Figure 12, again allowing for a comparison with Figure 5. Similar to observations at lower Reynolds numbers, there is negligible vorticity detachment from the cylinder. The primary tube displays analogous deformation patterns followed by wave propagation, consistent with the findings at lower Reynolds number. At later stages, increased distortion and emergence of smaller scales can be seen which are due to the increased tendency towards instability at this $Re_{\Gamma}$ value. Despite these developments, the tube survives the cut. The remnants left behind be observed in the last panel, at the end of the interaction. However, this tube has only about $85\%$ of the initial circulation left, as a fraction of it will have been lost to the small scale structures which arise and dissipate.
Figure 12: Visualization of the time evolution of $|\omega|$ for a simulation in the strong vortex regime with $Re_{\Gamma}=3000$, $I_{P}=0.25$. Snapshots are at $t\Gamma/D^{2}=200$, $240$, $280$, $320$ and $440$.
Additionally, we also show the vorticity magnitude through the $y=0$ plane just before the cylinder’s impact in the weak vortex regime for the two Reynolds numbers simulated. In this regime, the two cases look practically identical, aside from the higher values of vorticity and the thinner boundary layers. Notably, this shows that flow models which neglect secondary vorticity could adequately capture the interaction within this regime, primarily driven by deformations induced by the moving object.
Figure 13: Vorticity magnitude in the $z=0$ plane just before impact ($t\Gamma/D^{2}=220$) for $I_{P}=0.25$ and $Re_{\Gamma}=1000$ (left) and $Re_{\Gamma}=3000$ (right).
From these findings, it becomes evident that the primary control parameter in BVI is the impact parameter $I_{P}$. Varying the Reynolds number results in amplified distortion along the primary vortex, driven either by secondary vorticity or wave propagation. However, these alterations do not significantly impact the fundamental collision features, at least up to $Re_{\Gamma}=3000$. Furthermore, our simulations successfully reproduce and provide reasoning for most experimental outcomes, demonstrating that even in the absence of turbulence, the phenomenology of weak and strong BVI can be faithfully replicated.
@@ IV.1 Initial Remarks
We now move to an area which is much less analyzed in BVI: measurement of the force on the cylinder. However, an adequate definition of the force on is not simple. Two things have to be taken into account: first, the force coming from BVI is a deviation from a constant drag which resists the movement of the cylinder at a constant velocity in a quiescent flow. Secondly, this deviation will be localized to the point of interaction, and if we add up the forces for a quasi-infinite cylinder or blade undergoing BVI, they would be approximately the same for those where there is no BVI, as the constant drag would overwhelm the forces from BVI.
@@
To elaborate on this, in the left panel of Figure 14 we show the force coefficient distribution $C_{f,x}=f_{x}/(\frac{1}{2}\rho[\Gamma/\sigma]^{2})$ on the cylinder for the weak vortex case ($I_{P}=0.25$, $Re_{\Gamma}=1000$) before the cylinder has reached the vortex core. We note that this representation is highly distorted as the cylinder is very long. However, it is useful to note a few things. First, we can observe that the distribution is symmetric around $\theta=0$, which corresponds to the leading edge of the cylinder. This is a natural product of the symmetry of the system. Second, we note that while the region of strongest interaction is closely localized to the impact point, as could be expected, the return to asymptotic values is rather slow. Finally, we note that a rather small region of the cylinder sees an overall force in the direction of motion of motion of the cylinder ($C_{f,x}>0$). This is due to the fact that locally the flow direction can be reversed when the velocity induced due to the vortex is larger than the cylinder motion.
@@
We emphasize these results on the middle panel of Figure 14, which shows the force on the cylinder along its axis $C_{f}(z)$. The baseline case where no vortex is present is shown as a solid line. The forces can be seen to be somewhat symmetrical around $z=0$, i.e. the collision point. The overall force again shows positive values, coincident with regions where the velocity the cylinder sees is reversed. Finally, a long tail decay is observed: the forces do not return to the baseline case even at the edges of the cylinder. We note that our simulations yield qualitatively similar curves to those presented in the simplified models of Ref. [23], with the presence of a suction peaks and regions of increased pressure. However, comparison is difficult due to the simplifying assumptions made of the flow and the different parameter regimes investigated.
@@
Figure 14: Left: Pressure coefficient along the cylinder for $Re=1000$, $I_{P}=0.25$ at time $t\Gamma/D^{2}=160$. Middle: Force coefficient along the cylinder axis for the same case. Left: Evolution of drag coefficient with time for the full cylinder (orange) and half the cylinder (blue).
@@
As a final step we can calculate the average drag coefficient. If we use the entire extent of the cylinder, we obtain a value of $C_{F}=0.449$. If we integrate the force using the half of the cylinder closest to the vortex, and discard the two quarters close to the edges, we obtain a value of $C_{F}=0.428$. While both values are almost indistinguishable from the baseline value of $C_{F}=0.436$, one deviates upwards and the other downwards. These problems of adequately calculating $C_{F}$ are showcased in the right panel of Figure 14, which shows the time evolution of $C_{F}$ calculated using the full cylinder and half the cylinder. Both values show similar trends of increasing and decreasing during stages of the collision. However, they can be either above or below the baseline value depending on the cylinder segment considered for calculation. In any case, the deviations are mild, and only reach a maximum of ${}20\%$ drag increase for the half-cylinder calculation. We also note that some small oscillations are present in the force calculation, which have a frequency close to the natural frequency of the cylinder $D/V$.
@@ IV.2 Results
In this section, we show $C_{f}(z)$ for a few selected time instants for four characteristic cases. These are two cases for each of the strong ($I_{P}=0.05$) and weak ($I_{P}=0.25$) vortex regimes, at the two Reynolds numbers simulated. These are shown in Figure 15.
@@
Figure 15: Force coefficient along the cylinder axis at several time instances. Left column is cases at $I_{P}=0.05$, while right column is cases at $I_{P}=0.25$. Top row is at $Re_{\Gamma}=1000$, while bottom row is at $Re_{\Gamma}=3000$. Symbols: left column ($I_{P}=0.05$), light to dark: $t\Gamma/D^{2}=700$, $1100$, $1500$, $1900$ and $2300$. Right column ($I_{P}=0.25$), light to dark: $t\Gamma/D^{2}=160$, $200$, $220$, $240$, $280$ and $320$.
@@
We first start by discussing the case with $I_{P}=0.05$ and $Re_{\Gamma}=1000$. We notice that the curves are relatively similar to those shown in the previous subsection. The curves deviate from the baseline value, and they are roughly antisymmetrical around the collision point $z=0$. The deviations increase as the cylinder approaches the vortex, but they reach a maximum at distinct times: the peak force inversion ($C_{f}>0$) happens earlier than the peak suction ($C_{f}<0$). In the same manner, the values in the force inversion side return to the baseline values earlier than in the suction side. This is probably due to the fact that the cylinder boundary layer is being lifted up earlier from the force inversion side, so the interaction happens first on this half of the cylinder. Finally, we note that the deviations from the baseline value are much more significant in this case: after all, the vortex is strong. This means that it will create forces on the cylinder which are much larger than those originating solely from the ordinary drag force.
@@
Turning to the case with $I_{P}=0.25$ and $Re_{\Gamma}=1000$, we notice that the behaviour of the force has some similarities but also some differences when compared to the earlier case. The main difference is that the forces in the regions of large positive $C_{f}$ fluctuate much less- it attains a maximum value and remains close to that maximum value for most of the interaction. This suggests a probable absence of significant boundary layer liftoff. On the cylinder half experiencing force reversal, temporal variations are more pronounced yet localized due to the weaker vortex. In essence, little more can be extrapolated from these observations.
@@
Finally, we show the Reynolds number dependence of the force in the bottom panels of Figure 15, which can be contrasted to the top row. We note a relatively weak dependence of behavior on Reynolds number. The primary difference lies in the smaller baseline values of $C_{F}$, yet both curves exhibit analogous behaviour. Close to the impact point, spatial oscillations in $C_{f}(z)$ become more discernible, likely owing to the influence of smaller and more intense vortical structures on the cylinder.

### v2

@@ III.1 The distinction between weak and strong vortex regimes
Turning to the weak vortex regime, Figure 10 shows volume plots of the vorticity magnitude $|\omega|$ for $I_{P}=0.25$ and the two Reynolds numbers simulated at the early stages of the interaction. As in the strong vortex regime, two sources of vorticity are evident: the primary vortex and the wire’s boundary layer. However, the overall flow evolution looks very different when compared to Figure 5. The presence of a weaker vortex results in lower induced velocities on the wire. The wire’s boundary layer is only beginning to lift off in the left-most panels at $t\Gamma/D^{2}=200$ and $220$, and has no time to curl around and deform the primary vortex before impact. This can be appreciated by contrasting with the visualization at $t\Gamma/D^{2}=740$ in Figure 5 where the primary core is seen to have deformed on the right side.
Consequently, the wire’s boundary layer does not play a significant role in this scenario. This lack of interaction before the wire approaches the core serves as a key characteristic of the weak vortex regime, and has historically facilitated its modeling using inviscid methods. Instead, the physics in this regime are dominated by the displacement and deformation of the core, which is propagated through axial waves, as illustrated in the visualizations at $t\Gamma/D^{2}=240$ and onwards. The core is deformed due to the physical presence of the wire, which forces oppositely signed vorticity to approach the core in a process reminiscent of vortical reconnection [32] (marked by a blue arrow in Figure 10). This displaces the vortex tube, and stretches it into a thin “neck” until the core is eventually “cut”. This process is more marked in the $Re_{\Gamma}=3000$ case, where the vorticity is intensified substantially.
@@
Simultaneously, the core deformations close to the wire propagate through waves along the primary tube’s axis, as was previously observed in Figure 3. These waves do not substantially modify the radius of the primary vortex, a phenomena observed in the simulations and experiments of weak-vortex BVI with axial flow in Ref. [22]. Here, there is a crucial difference from the experiments with axial flow which introduces a notable asymmetry in wave propagation. Our simulations instead show a symmetrical propagation originating from the cylinder impact zone.
@@
Figure 10: Volume rendering visualization of the time evolution of $|\omega|$ during the early stages for a simulation in the weak vortex regime $I_{P}=0.25$ with $Re_{\Gamma}=1000$ (top row) and $Re_{\Gamma}=3000$ (bottom row).
As in the previous section, a closer look is provided by looking at two-dimensional cuts. Figure 11 represents the vorticity magnitude plotted at the central $x$-$y$ plane at three time instants as the wire is approaching the tube for the two Reynolds numbers, and Figure 12 shows the vorticity through the $x$-$z$ plane for the same time instants and the two Reynolds numbers.
@@
The Reynolds number dependence is first reflected as thinner and more intense boundary layers. For the first panel at $t\Gamma/D^{2}=220$, the secondary vorticity has not had time to completely wrap around the core, and is not visible in the $z=0$ cut. The primary vortex has been deformed by the presence of the wire which is pushing oppositely signed vorticity towards it. Notably, although the vorticity has less time to detach, it exhibits a higher value than in Figure 7. This can be attributed to the non-dimensionalization using $\Gamma/\sigma$ as a velocity scale, where higher values of $I_{P}$ lead to greater wire velocities, resulting in larger vorticity values for a boundary layer when normalized using $\Gamma$-$\sigma$ units.
By $t\Gamma/D^{2}=240$, the secondary vorticity has had time to begin to wrap around the primary vortex. As suggested by Ref. [22], the primary vortex can entrain a small quantity of vorticity from the boundary layer during the interaction. However, it is almost indiscernible for the panel at $Re_{\Gamma}=1000$, and for $Re_{\Gamma}=3000$, while visible, it is eclipsed by the vorticity intensification due to the core stretching and deformation. The intensification of vorticity is particularly marked for the higher Reynolds number, a phenomena which is again reminiscent of reconnection, where increases in Reynolds number produce increasingly thinner vortex sheets of large vorticity magnitudes [42]. However, once viscosity takes over, annihilating or reconnecting the primary and opposite vorticity, this does not seem to have a bearing on the rest of the flow. This can explain why flow models which neglect secondary vorticity can adequately capture the interaction within this regime: the phenomenology is primarily driven by the deformations induced by the moving object, rather than the annihilation of vorticity.
By the latest time, $t\Gamma/D^{2}=260$, the cylinder is cutting the vortex, and almost no vorticity is visible in the plots close to the wire. By returning to the three-dimensional visualization of Figure 10, it is clear that the cut has taken place. In these simulations, some of the phenomena usually attributed to the bending regime can be observed, i.e. deformations of the primary vortex, and the stretching (or “necking”) of the core near the impact location. This hints at the fact that the transition between strong vortex and bending regimes, which is outside the scope of this paper, will also be smooth.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 11: Vorticity magnitude $|\omega|$ at the $z=0$ plane evolution just before impact in the weak vortex regime with $I_{P}=0.25$ at $Re_{\Gamma}=1000$ (top row) and $Re_{\Gamma}=3000$ (bottom row). Snapshots are at $t\Gamma/D^{2}=220$ (left), $240$ (middle) and $260$ (right). Abbreviations: “PVN” primary vortex necking, “DSV” detached secondary vorticity.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
Figure 12: Vorticity magnitude $|\omega|$ at the $y=0$ mid-plane evolution for a simulation in the weak vortex regime at $I_{P}=0.25$ with $Re_{\Gamma}=1000$ (top row) and with $Re_{\Gamma}=3000$ (bottom row). Snapshots are at $t\Gamma/D^{2}=220$ (left), $240$ (middle), $260$ (right). Abbreviations: “DSV” detached secondary vorticity, “OV” opposite vorticity, “PV” primary vortex.
The flow phenomenology after the wire has completely transversed the vortex is shown in full three dimensional visualizations in Figure 13, and in a selected cut in the $y=0$ plane in Figure 14. As the wire passes through the vortex, it leaves behind a wake. The waves reach the limit of the computational domain around $t\Gamma/D^{2}=400$. The primary vortex reconstitutes itself, even if vortex loops are seen to detach from it, especially at high Reynolds number. The resulting distribution of vorticity is different from the original Gaussian, as it is now concentrated in a thin sheet and displaced to the right by a few vortex cores. Despite this, a comparison of the initial and final circulation of the vortex for $Re_{\Gamma}=1000$ reveals only a $5\%$ loss throughout the entire process, while at $Re_{\Gamma}=3000$ the core has lost $15\%$ of its original circulation, very likely to the small scale structures which arise and dissipate. This showcases the vortex’s ability to withstand the interaction significantly better than the case at $I_{P}=0.05$, which suffered a $75\%$ loss of its original circulation at $Re_{\Gamma}=1000$, and was not significantly detectable at $Re_{\Gamma}=3000$.
Figure 13: Volume visualization of the time evolution of $|\omega|$ during the late stages for a simulation in the weak vortex regime $I_{P}=0.25$ with $Re_{\Gamma}=1000$ (top row) and $Re_{\Gamma}=3000$ (bottom row).
Refer to captionRefer to caption
Figure 14: Vorticity magnitude $|\omega|$ at the $y=0$ mid-plane evolution and $t\Gamma/D^{2}=400$ for a simulation in the weak vortex regime at $I_{P}=0.25$ with $Re_{\Gamma}=1000$ (left) and with $Re_{\Gamma}=3000$ (right).
@@
The loss of circulation after the interaction has taken place ($\Gamma_{f}$) can be quantified for all cases at $Re_{\Gamma}=1000$. Figure 15 shows $\Gamma_{f}/\Gamma_{0}$, $\Gamma_{0}$ being the original circulation of the vortex. The amount of circulation in the vortex by the end of the interaction monotonically increases with $I_{P}$ in the range studied. We note that this trend cannot persist, as the circulation must saturate at $\Gamma_{f}\approx\Gamma_{0}$ at higher values of $I_{P}$, as it is not expected that the wire imparts additional circulation to the vortex.
@@
From these findings, it becomes evident that the primary control parameter in BVI is the impact parameter $I_{P}$. Varying the Reynolds number results in vorticity intensification and amplified distortion along the primary vortex, driven either by secondary vorticity or wave propagation. Also, at the later stages finer flow structures and turbulent clouds appear at higher Reynolds number, due to the increased instability of the vortices. However, these alterations do not significantly impact the fundamental collision features, at least up to $Re_{\Gamma}=3000$. Furthermore, our simulations successfully reproduce and provide reasoning for most experimental outcomes, demonstrating that even in the absence of turbulence, the phenomenology of weak and strong BVI can be faithfully replicated.
@@ IV.1 Initial Remarks
This section discusses an area which is much less analyzed in BVI: measurement of the force on the cylinder. However, an adequate definition of the force on is not simple. Two things have to be taken into account: first, the force coming from BVI is a deviation from a constant drag which resists the movement of the cylinder at a constant velocity in a quiescent flow. Secondly, this deviation will be localized to the point of interaction, and if the forces for a quasi-infinite cylinder or blade undergoing BVI are added up, they would be approximately the same for those where there is no BVI, as the constant drag would overwhelm the forces from BVI.
@@
To elaborate on this, the left panel of Figure 16 shows the force coefficient distribution $C_{f,x}=f_{x}/(\frac{1}{2}\rho[\Gamma/\sigma]^{2})$ on the cylinder for the weak vortex case ($I_{P}=0.25$, $Re_{\Gamma}=1000$) before the cylinder has reached the vortex core. We note that this representation is highly distorted as the cylinder is very long. However, it is useful to note a few things. First, the distribution is symmetric around $\theta=0$, which corresponds to the leading edge of the cylinder. This is a natural product of the symmetry of the system. Second, while the region of strongest interaction is closely localized to the impact point, as could be expected, the return to asymptotic values is rather slow. Finally, a rather small region of the cylinder sees an overall force in the direction of motion of motion of the cylinder ($C_{f,x}>0$). This is due to the fact that locally the flow direction can be reversed when the velocity induced due to the vortex is larger than the cylinder motion.
@@
These results are emphasized in the middle panel of Figure 16, which shows the force on the cylinder along its axis $C_{f}(z)$. The baseline case where no vortex is present is shown as a solid line. The forces can be seen to be somewhat symmetrical around $z=0$, i.e. the collision point. The overall force again shows positive values, coincident with regions where the velocity the cylinder sees is reversed. Finally, a long tail decay is observed: the forces do not return to the baseline case even at the edges of the cylinder. We note that the simulations yield qualitatively similar curves to those presented in the simplified models of Ref. [23], with the presence of a suction peaks and regions of increased pressure. However, comparison is difficult due to the simplifying assumptions made of the flow and the different parameter regimes investigated.
Refer to captionRefer to captionRefer to caption
@@
Figure 16: Left: Pressure coefficient along the cylinder for $Re_{\Gamma}=1000$, $I_{P}=0.25$ at time $t\Gamma/D^{2}=160$. Middle: Force coefficient along the cylinder axis for the same case. Left: Evolution of drag coefficient with time for the full cylinder (orange) and half the cylinder (blue).
@@
As a final step, the average drag coefficient can be calculated. If the entire extent of the cylinder is used, a value of $C_{F}=0.449$ is obtained. If the cylinder half closest to the vortex is used, and the two quarters close to the edges are discarded, a value of $C_{F}=0.428$ is obtained. While both values are almost indistinguishable from the baseline value of $C_{F}=0.436$, one deviates upwards and the other downwards. These problems of adequately calculating $C_{F}$ are showcased in the right panel of Figure 16, which shows the time evolution of $C_{F}$ calculated using the full cylinder and half the cylinder. Both values show similar trends of increasing and decreasing during stages of the collision. However, they can be either above or below the baseline value depending on the cylinder segment considered for calculation. In any case, the deviations are mild, and only reach a maximum of ${}20\%$ drag increase for the half-cylinder calculation. We also note that some small oscillations are present in the force calculation, which have a frequency close to the natural frequency of the cylinder $D/V$.
@@ IV.2 Results
In this section, $C_{f}(z)$ is calculated for a few selected time instants for four characteristic cases. These are two cases for each of the strong ($I_{P}=0.05$) and weak ($I_{P}=0.25$) vortex regimes, at the two Reynolds numbers simulated. These are shown in Figure 17.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 17: Force coefficient along the cylinder axis at several time instances. Left column is cases at $I_{P}=0.05$, while right column is cases at $I_{P}=0.25$. Top row is at $Re_{\Gamma}=1000$, while bottom row is at $Re_{\Gamma}=3000$. Symbols: left column ($I_{P}=0.05$), light to dark: $t\Gamma/D^{2}=700$ (light red), $1100$ (dark red), $1500$ (black), $1900$ (dark blue) and $2300$ (blue). Right column ($I_{P}=0.25$): $t\Gamma/D^{2}=160$ (light red), $200$ (dark red), $240$ (black), $280$ (dark blue) and $320$ (light blue). The horizontal thin line represents the baseline value.
@@
The discussion starts with the case with $I_{P}=0.05$ and $Re_{\Gamma}=1000$. The curves are relatively similar to those shown in the previous subsection. The curves deviate from the baseline value, and they are roughly antisymmetrical around the collision point $z=0$. The deviations increase as the cylinder approaches the vortex, but they reach a maximum at distinct times: the peak force inversion ($C_{f}>0$) happens earlier than the peak suction ($C_{f}<0$). In the same manner, the values in the force inversion side return to the baseline values earlier than in the suction side. This is probably due to the fact that the cylinder boundary layer is being lifted up earlier from the force inversion side, so the interaction happens first on this half of the cylinder. Finally, the deviations from the baseline value are much more significant in this case: after all, the vortex is strong. This means that it will create forces on the cylinder which are much larger than those originating solely from the ordinary drag force.
@@
Turning to the case with $I_{P}=0.25$ and $Re_{\Gamma}=1000$, the behaviour of the force has some similarities but also some differences when compared to the earlier case. The main difference is that the forces in the regions of large positive $C_{f}$ fluctuate much less- it attains a maximum value and remains close to that maximum value for most of the interaction. This suggests a probable absence of significant boundary layer liftoff. On the cylinder half experiencing force reversal, temporal variations are more pronounced yet localized due to the weaker vortex. In essence, little more can be extrapolated from these observations.
@@
Finally, the bottom panels of Figure 17 show the Reynolds number dependence of the force, which can be contrasted to the top row. There is a relatively weak dependence of behavior on Reynolds number. The primary difference lies in the smaller baseline values of $C_{F}$, yet both curves exhibit analogous behaviour. Close to the impact point, spatial oscillations in $C_{f}(z)$ become more discernible, likely owing to the influence of smaller and more intense vortical structures on the cylinder.